E I A Study 36 us showed twitter 7 & College Students FB. 9o at least one FB
Answer #1

ml FB) = ns. nl twilkes / = 36 nlat leest one) = n(FD) + N(T)-n (bath) - go = 25 + 36 - - - n = 11-9. = 21 Scanned With I Cam